OpenStudy (anonymous):

Does the following have a solution: {y=2x+3 {y=3x+2

OpenStudy (anonymous):

Yes. 2x+3=3x+2 3=x+2 x=1 Then plug in your x value in either of your other equations. y=2(1)+3 y=2+3 y=5 To check you can plug both of them in to either of your original equations. 5=2(1)+3 5=2+3 5=5